### v1

# BandRC: Band Shifted Raised Cosine Activated Implicit Neural Representations

Abstract
In recent years, implicit neural representations(INRs) have gained popularity in the computer vision community. This is mainly due to the strong performance of INRs in many computer vision tasks. These networks can extract a continuous signal representation given a discrete signal representation. In previous studies, it has been repeatedly shown that INR performance has a strong correlation with the activation functions used in its multilayer perceptrons. Although numerous activation functions have been proposed that are competitive with one another, they share some common set of challenges such as spectral bias(Lack of sensitivity to high-frequency content in signals), limited robustness to signal noise and difficulties in simultaneous capturing both local and global features. and furthermore, the requirement for manual parameter tuning. To address these issues, we introduce a novel activation function, Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) tailored to enhance signal representation capacity further. We also incorporate deep prior knowledge extracted from the signal to adjust the activation functions through a task-specific model. Through a mathematical analysis and a series of experiments which include image reconstruction (with a +8.93 dB PSNR improvement over the nearest counterpart), denoising (with a +0.46 dB increase in PSNR), super-resolution (with a +1.03 dB improvement over the nearest State-Of-The-Art (SOTA) method for 6X super-resolution), inpainting, and 3D shape reconstruction we demonstrate the dominance of BandRC over existing state of the art activation functions.
1 Introduction
With newly emerging Implicit Neural Networks (INRs), signal representation has taken a further step. INRs are a novel type of neural network capable of learning continuous representations of discrete signals. INRs have shown ‘remarkable performance in representing a variety of signals such as images, audio, video, 3D objects [1, 5, 6] and even scenes  [24, 10]. Traditional explicit signal representation methods store signal values discretely on coordinate grids. Although they achieve sufficient results, they struggle when dealing with high-dimensional data. Particularly, computational cost and memory capacity rise exponentially with dimensionality and resolution of data [14]. For example, in 3D object representation voxels are used to store the discrete containers. The memory footprint of voxel representations grows cubically with resolution. However, unlike these traditional methods, INRs take a new approach to training neural networks to approximate the relationship between input coordinates and their values using a continuous function.
Due to the strong ability of INRs to effectively learn and represent complex data patterns, they have been widely investigated in recent studies in the context of many computer vision tasks. They have demonstrated applications in several inverse problems, including image denoising, image inpainting, and super-resolution. Despite this, INR methods suffer from spectral bias  [16, 29] (learning bias towards the low frequency components) which hinders their ability to represent sharp details in signals. Traditional activation functions such as ReLU, Sigmoid, and Tanh heavily suffer from this spectral bias  [25, 26, 16]. To address this issue,  [26] proposed to pass the input coordinates through a simple Fourier feature mapping which enables multilayer perceptron (MLP) to learn high-frequency functions. This approach is known as positional embeddings in INR literature. Following this, a variety of activation functions are invented and they can be used with positional embeddings to further expand their frequency support. However, these activation functions still struggle when representing complex signals with fine details. In most of the INR methods, data noise is considered to be zero which is not the case in practical applications. Moreover, they still tend to struggle with more complex tasks such as super-resolution, image inpainting and 3D object reconstruction. Given that robust performance in such complex tasks is presumed to stem from the inherent capabilities of an INR, these limitations are a revelation of constraints in current implementations.
To address the aforementioned issues, we propose a novel activation function Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) which, due to its inherent nature is capable of grasping finer information from signals. In the frequency domain, the proposed activation function takes the form of a frequency-shifted raised cosine filter which offers controlled bandwidth and center frequency while offering frequency compactness. Not only that, but similar to WIRE [19], BandRC offers compactness in spatial domain. Furthermore, we leverage the method introduced by  [9] to obtain the optimal activation function parameters through a task-specific model. Through extensive mathematical and experimental analysis, we show that the BandRC activation function offers superior signal representation and outperforms current state-of-the-art activation functions in terms of higher representation capability and noise robustness. Specifically, we show that our activation function achieves superior performance not only in signal representation but also in other computer vision inverse problems such as denoising, super-resolution, image inpainting.
In summary, we make the following contributions
  1. 1.
    Introduce a novel activation function with better frequency support that mitigates spectral bias and surpasses State-Of-The-Art activation(SOTA) functions.
  2. 2.
    Demonstrate, through mathematical analysis and experimental procedures, the resilience to spectral bias of the proposed model.
  3. 3.
    Incorporate the leveraging prior knowledge through a task specific model [9] for the automated determination of activation function parameters.
  4. 4.
    Demonstrate the dominance of BandRC over other SOTA activation functions through a series of benchmark tests including signal representation and computer vision inverse tasks such as denoising, super-resolution, image inpainting.
2 Related Work
Implicit representations. According to the survey by  [3], INR research can be classified into three main parts, which are, enhancing position encoding, improving activation functions, advances in overall network architecture. Position encoding is the process of embedding high frequency content of the signal into the input coordinates of the INR to help them learn high-frequency components effectively. This approach can be used with many activation functions to achieve better performance [22]. Activation functions are specially formulated to increase the spectral support throughout the network. This is important in addressing the spectral-bias issue of INRs. In earlier implicit representations, Hanin and Rolnick  [7] have proposed ReLU-MLPs. While promising, the piecewise-linear nature of the ReLU function limits their ability to capture fine details and their ability to represent derivatives of the target signal. This is further confirmed using theoretical and experimental methods by Rahaman et al.  [16]. To overcome such issues, Sitzmann et al.  [25] proposed periodic activation functions (SIREN) where they use a sinusoid as an activation function. In SIREN, the activation function was defined as, $\phi_{(x)}=sin(\omega_{0}x)$. This activation function stands out due to the ability to model a broad spectrum of frequencies and its continuous nature, which allowed the gradients to be preserved. However, Saragadam et al. [19] showed that SIREN based models results in global artifacts when representing signals. They introduce a new wavelet based activation function -(WIRE), that can capture more details in the signal. Following this, many other activations were introduced. Some notable mentions are,Gaussian activations [17], Sinc [20], HOSC [21] and FINER [11].
Another key aspect of enhancing INR performance is the network architecture. A key weakness in INR models is the sensitivity of activation function parameters for each task. To address this issue, INCODE [9] proposes a novel model architecture that uses deep prior information of the signal to adjust the parameters of the activation function. This leads to better overall performance and most importantly, makes the INR model adaptable to the signal. Other significant approaches to improve model architecture include Fourier parameterized training [23].
Prior knowledge embedding and Conditional Neural Networks. In conditional neural networks, additional information and knowledge is taken into account during the learning phase. This additional information helps the network to adapt its output based on the given inputs. This is essentially embedding contextual information that depends on the task into the neural network. In the context of INRs, integrating latent codes derived from encoders has proven effective for representing complex high-dimensional data.  [18, 15, 13]. The approach proposed by  [9] makes use of a task-specific model tailored to each task together with a harmonizer network to provide contextual information into the INR. We leverage this method to dynamically adjust activation function parameters to guide them to their optimal potential based on the input signal.
3 METHOD
In this section, we will first introduce the basic definition of an implicit neural network. Then we will proceed to introduce the BandRC in detail.
Consider a signal sequence $\hat{S}_{x}$, sampled from a continuous signal $S_{x}:\mathbb{R}^{r}\rightarrow\mathbb{R}^{c}$. An implicit neural network is formulated using such a sampled data space to learn the essence of $S_{x}$. Such a neural network can be expressed as, $f_{\theta(x)}:\mathbb{R}^{r}\rightarrow\mathbb{R}^{c}$ where $\theta$ denotes the weights and biases of the network. The network is trained by minimizing the mean square error which is given by,
$L=\mathop{\mathbb{E}}_{x\in X}\|f_{\theta(x)}-\hat{S}_{x}\|^{2}$ (1)
The activation function of $f_{\theta(x)}$ plays a vital role in capturing the signal spectrum.
3.1 BandRC: Implicit Neural Representations with Raised Cosine Activations
To overcome the spectral bias and retain sharp information in input signals, we propose BandRC, a novel approach that can capture high-frequency components, and concurrently preserve local and global features better than previously introduced activation functions produces robust results. The architecture of BandRC INR is shown in Figure  1
Figure 1: Model Architecture.
Let $f_{\theta(x)}$ be defined as with $L$ hidden layers, each having neuron widths of $D_{1},D_{2},...,D_{L}$. Then $f_{\theta(x)}$ can be expressed as,
$f_{\theta(x)}=\phi_{L}(W^{L}z^{L-1}+b^{L})$ (2)
Where, $z^{i}$ denotes the pre-activation of the layer $L^{th}$ and $\phi_{L}$ denotes the activation function at $L^{th}$ layer. We propose setting the final layer with sigmoid activation (depending on the task) to bound the output to $[0,1]$. The rest of the L-1 layers are set with the proposed activation function. Now we will dive into the formulation of the activation function.
We aim to employ an activation such that its frequency response will take the form of a raised cosine filter. We further give it a frequency shift to have a better grasp at higher frequencies. The frequency response of our activation function is shown in Figure  2.
Figure 2: Frequency response of the proposed activation function
The proposed activation function can be written as the impulse response of this shifted raised cosine filter, which can be written as,
$\phi_{x}=\dfrac{1}{T}sinc\left(\dfrac{x}{T}\right)\dfrac{\cos\left(\dfrac{\pi\beta x}{T}\right)}{1-\left(\dfrac{2\beta x}{T}\right)^{2}}.\exp{(2\pi{\zeta}xj)}$ (3)
We suggest a fixed roll-off rate of $\beta=0.05$ and both bandwidth($T$) and the frequency shift($\zeta$) to be kept as learnable. The outputs at each layer are complex-valued, and the phase is preserved throughout the network. For a stable learning process, we propose to normalize activation function outputs and inputs at each layer and brim them into a unit circle. In the final layer, the real part of the output is extracted. The parameters of each activation function, $T$ and $\zeta$ depend on the training task and should be adjusted accordingly. For this, we leverage some prior knowledge extracted from the training data as suggested in [9]. This prior knowledge embedding process is done through a network block which is composed of a task-specific model followed by a harmonizer network. A latent code $z\in\mathbb{R}^{r}$ is generated from the pre-trained task-specific model and sent to the harmonizer network. The architecture of the harmonizer network was constructed following the approach in  [9], using linear layers followed by SiLU activations.
The harmonizer network then adjusts the activation function parameters in each training iteration. This process eliminates the need for initialization of the activation function parameters, offering a superior benefit over other conventional INR models such as  [19].
shifting parameter ($\zeta$) The frequency shifting component dynamically shifts the base band spectrum of the RC filter, allowing better alignment with high-frequency features allowing better encoding and reconstruction. The prior embeddings feed the $\zeta$ to adhere to the harmonizing atoms to preserve essential frequency components. In other terms the shift repositions signal energy in the Fourier domain. bandwidth parameter ($T$): The bandwidth component adapts the effective bandwidth for each activation. Influencing signal resolution without compromising smoothness and balancing aliasing effects, leading to improved flexibility. Roll-off factor being a constant ensures controlled transition regions, avoiding excessive oscillations. Collectively these independent parameters ensure the capability of the activation to better capture sharp edges, fine textures, and rapid variations in signals, making them more efficient for implicit continuous representations.
3.2 A mathematical perspective of the activation function
In this section, we will try to analyze the expressive power of the proposed activation function using theoretical foundations from  [29, 12]. According to  [29], using polynomial series expansion, any activation function can be written in the form,
$\rho(x)=\sum_{k=1}^{\infty}\alpha_{k}x^{k}$ (4)
Now consider a single frequency mapping $\gamma(r)=e^{j\omega r}$ going through such activation function $\rho(x)$. We can express this as [29],
$\rho(\gamma(r))=\sum_{k=1}^{\infty}\alpha_{k}e^{j\omega kr}$ (5)
Now according to  [29], this harmonic expansion controls the frequency representation in INRs. it can be observed that INR families are analogous to structured signal dictionaries whose atoms are integer harmonics of the set of initial mapping frequencies. Leveraging this observation, we further demonstrate the power of spectral representation capabilities using a polynomial approximation of BandRC activation. According to  [12] each nonlinear layer creates frequency harmonics, whose magnitude increases with the network depth. They call it the ”Blue-shift”, and this provides a global view of how network architectures behave.
Stone-Weierstrass theorem further says that it is possible to approximate any continuous function by a polynomial. Based on this, we approximate Raised Cosine, Gauss and Sin nonlinearities using Chebyshev polynomials of the first kind under the interval of [-5,5]. We compare the coefficient distribution up to 50 degrees as shown in  3.
Figure 3: Chebyshev coefficients of Raised Cosine, Sinusoid, Gauss
It can be observed that the Raised Cosine function has better stability over a greater number (up to 20) of polynomial coefficients(harmonic integers) comparatively. Both Sinusoidal and Gauss approximations coefficients degrade faster than Raised Cosine. Therefore, reducing the spectral support over a wide range of higher-frequency harmonics.
BandRC can be interpreted as the product of a frequency-shifting parameter (complex exponential) and a Raised Cosine function, similar to how WIRE [19] does with a Gaussian expression. We demonstrate that sinusoidal- and Gaussian-based activations exhibit poor stability compared to our proposed activation, making them less reliable in handling the spectral bias. Therefore, it can be concluded that the lower the degradation over the wider coefficient range, the greater the stability in representing frequency components of a signal. This makes BandRC ideal for mitigating spectral bias issues in signal representation.
3.3 Spectral Bias in INR: Experimental Setup
Now we will demonstrate the spectral bias of the proposed activation function using an experimental approach.
A 2D image of a chirp signal with a broader range of frequency components was used as the test signal. The spectral bias occurring during image reconstruction was calculated in the experiment. This was obtained by calculating the absolute error between the frequency spectrum of the ground truth and the frequency spectrum of the reconstructed image.
The experimental results of BandRC were compared with the results obtained from SIREN [25], WIRE [19] and INCODE [9]. According to Figure  4, BandRC shows less error (more black regions in spectral error plot)in the frequency domain compared to other SOTA methods. Thus, it experimentally proves that BandRC has the ability to learn both high and low-frequency components equally. Thus, mitigating the spectral bias problem.
Figure 4: Experimental setup for spectral bias demonstration
4 Experiments and Results
We utilize a 5-layered network shown in Fig.  1 equipped with 256 neurons 4 raised cosine-activated layers. Prior knowledge embedding model is specified for each task as discussed in  [9]. We perform our experiments in a Nvidia Quadro GV100 GPU with 32 GB memory. All the codes are written using PyTorch frameworks. We use Adam optimizer throughout all the experiments. Learning rates and scheduler parameters are given in each relevant section. We compare ore results with SIREN [25], WIRE [19], ReLU [7], INCODE [9]. More experimental details are discussed in the provided supplementary material.
4.1 Image Reconstruction
Implicit neural representations can be used to learn the implicit representation of a given image. This task is commonly referred as image reconstruction/representation. The INR network is given the input coordinates of an image and it is trained to predict the RGB values at each pixel.
Data. We conduct our experiments using the Kodak [4] dataset at native resolution. The reconstruction task is evaluated using the PSNR metric by comparing the reconstructed image with the ground-truth image.
Architecture. Similar to [9], we adopt a pre-trained ResNet34 [8] (trained on ImageNet1K_V1 [2]) truncated at the fifth layer as our task-specific model. The output features are globally average-pooled and passed to an MLP, akin to that in INCODE, to derive activation parameters, thereby adapting the activations to the input information bandwidth.
Observations. The image representation performance of BandRC compared with SOTA methods for Kodak [4] image 20 is presented in Fig. 5. It is apparent that BandRC clearly outperforms the nearest counterpart, INCODE by +8.93 dB and improves upon WIRE and SIREN by +12.2 dB and +12.63 dB respectively. In addition, BandRC shows clear improvements in clarity of the aircraft wheel, particularly compared to WIRE and SIREN.
Figure 5: Image representation of BandRC compared with SOTA methods.
4.2 Image Denoising
Data. In this section, we evaluate the robustness of INRs for noisy signals. DIV2K image dataset [27] was used in the following experiments. We add photon noise to the ground truth, where independent Poisson random variables are applied to each pixel. We set the mean photon count to 30 and the readout counts to 2 for all the image-denoising tasks.
Architecture. We implore the exact strategy used in image reconstruction to manipulate the activation parameters for the denoising task. We train the model with a learning rate of 0.005 and a decay rate of 0.1.
Observations. According to the experimental results for denoising, as shown in Fig. 6, BandRC has surpassed the SOTA methods. BandRC is 0.46 dB higher in the PSNR value compared to the second-highest INCODE. Further, from visual observations, BandRC has preserved the image colors better than SOTA methods. The perseverance of structural integrity can be highlighted in BandRc.
Figure 6: Image denoising of BandRC compared with SOTA methods.
4.3 Image Super-resolution
Image super-resolution is a classic machine learning task where the goal is to increase the resolution of the image. Often image is up-scaled by a factor of 4x or more while maintaining its details. This results in a high-resolution image of the original image. Because INRs are able to capture the inherent properties of a given image, a key advantage of INR models is that they can be used in image super-resolution tasks. In this section, we will evaluate the super-resolution capability of the proposed raised cosine activated INR model.
Data. An image from DIV2K image dataset [27] was used in the image super-resolution task. We downsample the original image by factors of 1/2 , 1/4 and 1/6 to use as ground truth to train the INR model. Then the 2x, 4x and 6x super-resolution tasks were performed and compared with the original image.
Architecture. Similar to the implementation of reconstruction and denoising. The model exploits the interpolation capabilities of the model.
Observations. The results are presented in Table 1 and it can be clearly seen that our method surpasses state-of-the-art methods by a considerable margin. In the visual representation of 6x super-resolution shown in Fig 7, one can observe the superiority of BandRC over other methods. Our method Gives the sharpest results while maintaining the original color contrast of the image. All the methods suffer from background noise. This can be clearly seen in the INCODE and SIREN methods, while BandRC gives the most accurate background reconstruction. Furthermore, in the enlarged wing section of the butterfly, our method preserves black colors well while resulting in the sharpest details compared to others while preserving the smoothness.
Method 2x PSNR 4x PSNR 6x PSNR
ReLU+PE 32.80 28.89 26.29
SIREN 32.26 29.62 27.31
WIRE 29.02 27.16 25.35
INCODE 32.83 29.96 26.63
BandRC 34.03 30.42 27.66
Table 1: BandRC vs. SOTAs in image super-resolution.
Figure 7: 6X Single image super-resolution - BandRC vs SOTA methods.
4.4 Image Inpainting
Image inpainting refers to the task of reconstructing missing or corrupted pixels of an image leveraging the surrounding pixel information. This task is specified in applications like, image restoration, artifact removal, and defect correction. Recent advancements in INR have exploited the potential of implicit continuous functions in enhancing image inpainting quality that is on par with the current theoretical approaches. This section we explore our competence with the existing SOTA models.
Data. Here, we utilize the Celtic spiral knots image with a resolution of 572x582x3. We sample 20 percent of the pixels from the original image to complete the reconstruction.
Architecture. We reconstruct the structural formation of the sampled pixels by applying a mask on the missing data. Then we feed the resnet34 with the reconstructed image to inherit the structural sense for the activation parameters. Finally, we inference the whole coordinate base to the model for the observations.
Observations. Fig. 8 depicts the performance of BandRC activation which surpasses the other modalities. The capability of BandRC has been able to secure a better PSNR ratio while maintaining a competitive SSIM score. This proves the continuity and the ability to recognize patterns from sparse signals. of the implicit function defined by BandRC.
Figure 8: Image inpainting of BandRC compared with SOTA methods.
4.5 Occupancy Fields
Implicit Neural reconstruction for 3D shapes have been able to encode traditional discrete grid-like coordinate structures to information-rich implicit continuous functions. Thus enabling smooth representations for higher resolutions. The signed distance functions (SDF) provide a negative value for points inside the surface, a positive value for points outside, and zero at the surface. Occupancy fields, on the other hand, provide a binary representation, marking points inside the object as occupied (1) and outside as empty (0). For 3D reconstruction tasks both information forms can be used to train implicit functions.
Data. For this task, we train the Thai statue for SOTA model comparison. First, we create an occupancy volume through point sampling on a 512x512x512 grid, assigning one and zeros to voxels within the object and outside the object respectively.
Architecture. Here, we use resnet3d-18  [28] as the task-specific model truncated up to layer 3. To infuse inherent local features of the whole 3D shape to acquire good the activation parameters appropriately. Thereby handling frequency shifting and filter bandwidth requirements. This enhances better feature representative qualities even in the complex edge-like structures on par with the current state-of-the-art activations.
Observations. The results, as presented in Fig. 9 and the Table 2 demonstrate the effectiveness of BandRC in occupancy fields representation. The results for other methods (WIRE, SIREN, ReLU+PE Intersection over Union (IOU) values) are obtained from  [19].
Figure 9: Occupancy fields representation of BandRC compared with SOTA methods
Method IOU
ReLU+PE 0.980
SIREN 0.970
WIRE 0.99
INCODE 0.991
BandRC 0.992
Table 2: BandRC vs. SOTAs in image occupancy fields.
5 Conclusion
We have proposed BandRC, an innovative activation function for implicit neural representations that achieves a higher-quality signal representation and less spectral bias. The proposed activation function takes the form of an impulse response of a shifted raised cosine filter. We explain the strong spectral support of BandRC through a new mathematical approach by utilizing Chebyshev polynomial approximations and through a detailed experimental setup. Through multiple benchmark experiments, we show that BandRC beats current state-of-the-art activation functions by a considerable margin in terms of capturing signal content, and solving inverse problems while being robust to signal noise and less sensitive to parameter variations.

### v2

@@
 We also incorporate deep prior knowledge extracted from the signal to adjust the activation functions through a task-specific model.
-Through a mathematical analysis and a series of experiments which include image reconstruction (with a +8.93 dB PSNR improvement over the nearest counterpart), denoising (with a +0.46 dB increase in PSNR), super-resolution (with a +1.03 dB improvement over the nearest State-Of-The-Art (SOTA) method for 6X super-resolution), inpainting, and 3D shape reconstruction we demonstrate the dominance of BandRC over existing state of the art activation functions.
+Through a mathematical analysis and a series of experiments which include image reconstruction (with an average PSNR improvement of +5.67 dB over the nearest counterpart across a diverse image dataset), denoising (with a +0.46 dB increase in PSNR), super-resolution (with a +1.03 dB improvement over the nearest State-Of-The-Art (SOTA) method for 6X super-resolution), inpainting, and 3D shape reconstruction we demonstrate the dominance of BandRC over existing state of the art activation functions.
 1 Introduction
@@
 Not only that, but similar to WIRE [19], BandRC offers compactness in spatial domain.
-Furthermore, we leverage the method introduced by  [9] to obtain the optimal activation function parameters through a task-specific model.
+Furthermore, we leverage the method introduced by  [9] to obtain the optimal activation function parameters through a prior-knowledge embedding model.
+The output of this prior-knowledge model is fed to a special regularization mechanism to boost the optimization process and further enhance the results.
 Through extensive mathematical and experimental analysis, we show that the BandRC activation function offers superior signal representation and outperforms current state-of-the-art activation functions in terms of higher representation capability and noise robustness.
@@
   3. 3.
-    Incorporate the leveraging prior knowledge through a task specific model [9] for the automated determination of activation function parameters.
+    Incorporate the leveraging of prior knowledge through a task-specific model [9] for the automated determination of activation function parameters, together with a specialized regularization mechanism.
   4. 4.
@@
 The outputs at each layer are complex-valued, and the phase is preserved throughout the network.
-For a stable learning process, we propose to normalize activation function outputs and inputs at each layer and brim them into a unit circle.
+For a stable learning process, we propose to normalize activation function outputs and inputs at each layer and bring them into a unit circle on the complex plane.
 In the final layer, the real part of the output is extracted.
@@
 The architecture of the harmonizer network was constructed following the approach in  [9], using linear layers followed by SiLU activations.
-The harmonizer network then adjusts the activation function parameters in each training iteration.
-This process eliminates the need for initialization of the activation function parameters, offering a superior benefit over other conventional INR models such as  [19].
-shifting parameter ($\zeta$) The frequency shifting component dynamically shifts the base band spectrum of the RC filter, allowing better alignment with high-frequency features allowing better encoding and reconstruction.
+The harmonizer network then sends its output to special regularizer blocks attached to each activation function.
+The regularization mechanism uses a sigmoid projection within some pre-specified bounds 4.
+$\theta=a+(b-a).sigmoid(\hat{\theta})$ (4)
+Here, a and b are the lower and upper bounds for the activation function parameters($T,\zeta$).
+They should be pre-specified before training.
+This harmonizer network, together with the regularization mechanism, then adjusts the activation function parameters in each training iteration.
+This process eliminates the need for initialization of the activation function parameters, and boosts overall performance, offering a superior benefit over other conventional INR models such as  [19].
+The frequency shifting component($\zeta$) dynamically shifts the base band spectrum of the RC filter, allowing better alignment with high-frequency features allowing better encoding and reconstruction.
 The prior embeddings feed the $\zeta$ to adhere to the harmonizing atoms to preserve essential frequency components.
 In other terms the shift repositions signal energy in the Fourier domain.
-bandwidth parameter ($T$): The bandwidth component adapts the effective bandwidth for each activation.
+The bandwidth component($T$) adapts the effective bandwidth for each activation.
 Influencing signal resolution without compromising smoothness and balancing aliasing effects, leading to improved flexibility.
-Roll-off factor being a constant ensures controlled transition regions, avoiding excessive oscillations.
-Collectively these independent parameters ensure the capability of the activation to better capture sharp edges, fine textures, and rapid variations in signals, making them more efficient for implicit continuous representations.
+Roll-off factor($\beta$) being a constant ensures controlled transition regions, avoiding excessive oscillations.
+Collectively, these independent parameters ensure the capability of the activation to better capture sharp edges, fine textures, and rapid variations in signals, making them more efficient for implicit continuous representations.
 3.2 A mathematical perspective of the activation function
@@
 It can be observed that the Raised Cosine function has better stability over a greater number (up to 20) of polynomial coefficients(harmonic integers) comparatively.
-Both Sinusoidal and Gauss approximations coefficients degrade faster than Raised Cosine.
+Both Sinusoidal and Gauss approximated coefficients degrade faster than Raised Cosine.
 Therefore, reducing the spectral support over a wide range of higher-frequency harmonics.
@@
 Figure 4: Experimental setup for spectral bias demonstration
+3.4 Spectral Bias in INR: Layer-wise frequency analysis
+A layer-wise frequency analysis was done on BandRC and SIREN activations to experiment with the spectral bias further.
+A chirp image was used to experiment with the capability of representing high-frequency information as seen in Figure  5.
+The average magnitude spectrum of each layer output was obtained for the chirp image.
+The heights of the magnitude spectra along the x-axis of the 2-D Fourier transforms for layers 1-5 are shown in Figure  5.
+Accordingly, BandRC has higher magnitudes at high frequency components compared to SIREN.
+As mentioned in  [9], this confirms BandRC is capable in representing better high frequency information.
+Figure 5: Layer-wise frequency response for BandRC compared to SIREN
 4 Experiments and Results
@@
 The image representation performance of BandRC compared with SOTA methods for Kodak [4] image 20 is presented in Fig.
-5.
+7.
 It is apparent that BandRC clearly outperforms the nearest counterpart, INCODE by +8.93 dB and improves upon WIRE and SIREN by +12.2 dB and +12.63 dB respectively.
 In addition, BandRC shows clear improvements in clarity of the aircraft wheel, particularly compared to WIRE and SIREN.
+Furthermore, we present BandRC’s performance on 24 Kodak lossless images in Fig.
+6 (a), which demonstrates it clearly outperforms the SOTA activation functions.
+It is clear that BandRC consistently achieves the highest performance, averaging 41.24 dB PSNR, significantly exceeding Incode’s 35.57 dB.
+These results are shown in Table 1.
+Also, notably, BandRC converges faster compared to other activation functions while maintaining stability as shown in Fig.
+6 (b).
+Figure 6: BandRC performance analysis on image reconstruction
+Activation Average PSNR
+ReLU+PE 32.79
+SIREN 33.09
+WIRE 29.41
+INCODE 35.57
+BandRC 41.24
+Table 1: Average PSNR performance for ImagesBandRC vs. SOTAs in Kodak image reconstruction.
 Figure 7: Image representation of BandRC compared with SOTA methods.
@@
 According to the experimental results for denoising, as shown in Fig.
-6, BandRC has surpassed the SOTA methods.
+8, BandRC has surpassed the SOTA methods.
 BandRC is 0.46 dB higher in the PSNR value compared to the second-highest INCODE.
@@
 Occupancy fields, on the other hand, provide a binary representation, marking points inside the object as occupied (1) and outside as empty (0).
-For 3D reconstruction tasks both information forms can be used to train implicit functions.
+For 3D reconstruction tasks, both information forms can be used to train implicit functions.
 Data.
 For this task, we train the Thai statue for SOTA model comparison.
-First, we create an occupancy volume through point sampling on a 512x512x512 grid, assigning one and zeros to voxels within the object and outside the object respectively.
+First, we create an occupancy volume through point sampling on a 512x512x512 grid, assigning ones and zeros to voxels within the object and outside the object, respectively.
 Architecture.
 Here, we use resnet3d-18  [28] as the task-specific model truncated up to layer 3.
-To infuse inherent local features of the whole 3D shape to acquire good the activation parameters appropriately.
+To infuse the inherent local features of the whole 3D shape to acquire good activation parameters appropriately.
 Thereby handling frequency shifting and filter bandwidth requirements.
@@
 Table 3: BandRC vs. SOTAs in image occupancy fields.
+Table 4 reports results for, image representation task with Kodak 22 image, the peak signal-to-noise ratio (PSNR) performance deviations of the proposed BandRC activation for 1000 training epochs under varying architectural configurations.
+Empirical evidence suggests that the optimal learning rate is sensitive to the model’s depth, and has accordingly been adjusted for each configuration, as indicated in the table.
+Additionally, we observe a gradual increase in the maximum attainable PSNR with respect to the layer width, highlighting a strong correlation between the number of neurons and the representational capacity induced by the BandRC activation.
+Notably, the model achieves a PSNR of approximately 52 dB when configured with a width of 512 neurons and a depth of 4 hidden layers, which is an exceptional performance benchmark in this context.
+However, to ensure a balance between computational efficiency and reconstruction fidelity, we adopted a configuration with a width of 256 and a depth of 3 hidden layers.
+This setup yields a PSNR of 39.57 dB, which remains close to the 40 dB threshold and further substantiates the scalability and effectiveness of the proposed nonlinear activation function.
+Depth Width 64 128 256 512
+2 (lr = 0.01) 28.52 32.92 38.09 42.84
+3 (lr = 0.01) 34.40 35.35 39.57 47.80
+4 (lr = 0.001) 29.47 35.01 37.19 52.00
+Table 4: PSNR ablation study on width/depth in dB
 5 Conclusion